Find a standard form equation for the line through the pair of points.
(-3,-8) and (4,-1)

Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(16877)



Use the two-point form of an equation of a line:



where and are the coordinates of the given points.

After plugging in the numbers as indicated, to the indicated arithmetic, then use the standard tools for manipulating equations to put your equation into standard form, namely:

Note, some texts/professors require that the coefficients in standard form be integers. Check with your instructor.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Alan3354(30993)

Determinants is the best method.
----------
|+y +y 1|
|-3 -8 1| = 0
|+4 -1 1|
-----------
x(-8+1) -y(-3-4) + 3+32 = 0
-7x + 7y + 35 = 0
x - y = 5